OpenStudy (anonymous):

The equation d = 2.5t describes the distance in meters d covered by a canoe-racing team in t seconds. How long does it take the team to go 125 meters? How long does it take them to go 400 meters? would i multiply 2.5 by 125 to find the first answer!!???

OpenStudy (anonymous):

d = 2.5t Part A. 125 meters is the distance; d= 125 125 = 2.5 t t = 125/2.5 t = 50 seconds Part B. 400 meters is the distance; d = 400 400 = 2.5t t = 400/2.5 t = 160 seconds, or 2 minutes and 40 seconds